# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

### v1

@@ 3.2. Analysis for Rule 1
((3.26)) $\Omega(1+\eta)k_{*}(y^{\delta})\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
Hence as a result, we get
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2},$
which completes the proof. ∎
@@
Let the assumptions of Lemma 3.1 be fulfilled. Moreover assume that IRSGD method (3.1) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
In particular, for the case $y^{\delta}=y^{\dagger}$ ((((i.e., if $\delta=0)$, we have
$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty,$
where $D_{1}$ is defined in (3.22).
@@
((3.27)) $D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
@@
((3.28)) $\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
Next, for the case $\delta=0$, it follows directly from (3.27) that
((3.29)) $\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
Here we have used the fact that $\sum_{l=0}^{\infty}\lambda_{k}<\infty.$ Thus, (3.28) and (3.29) complete the proof. ∎
@@
If $\lambda_{\text{max}}=0$, the parameter $\kappa$ in conditions (3.7) and (3.22) can be set to zero. This adjustment mirrors the conditions required to establish the same results for the stochastic gradient descent (SGD) method (1.9).
3.3. Convergence for non noisy case
The following result illustrate that the sequence of IRSGD iterates $\{u_{k}\}_{k\geq 1}$ in (3.1) with exact data $y$ forms a Cauchy sequence.
@@ 3.3. Convergence for non noisy case
Let all the assumptions of Lemma 3.1 hold. Additionally assume that
((3.30)) $D_{2}:=(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})\geq 0.$
Then the IRSGD method (3.1) for $\delta=0$ produces a sequence of iterates $\{u_{k}\}_{k\geq 1}$ which is a Cauchy sequence.
@@
For a solution $u^{\dagger}$ of (1.2), we define $e_{k}:=u_{k}-u^{\dagger}$. As indicated in part (b) of Lemma 3.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$. Let it converge to some $\epsilon\geq 0$. It follows directly from 3.2 that, for $\delta=0$ we have
@@
From the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we deduce that
@@
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
$-\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
By using (3.30) and considering the full conditional, we reach at
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+D_{2}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
@@
for $m<k,$ where we have used the formula
@@
These estimates, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, immediately imply
@@
Further, to show the convergence of the iterates $u_{k}$, it is crucial to note that $\mathbb{E}[\langle\cdot,\cdot\rangle]$ serves as an inner product. For any $n\geq k$, we select an index $m$ satisfying $n\geq m\geq k$ such that
((3.33)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\forall k\leq i\leq n.$
Using inequality $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]|^{1/2}$ and the identities
$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{j},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
$\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
it is sufficient to establish that $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ tend to zero as $k\rightarrow\infty$. As $k\rightarrow\infty$, the last two terms on the right-hand sides of the preceding two identities tend to $\epsilon-\epsilon=0$. This is due to the convergence of $\mathbb{E}[\|e_{k}\|^{2}]$ to $\epsilon$, as discussed above. Next, we proceed to demonstrate that the term $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also approaches to zero as $k\rightarrow\infty$. In fact, by virtue of the definition of $u_{k}$, we have the following representation
$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}).$
for $m<k$. Using this formula, we obtain the estimate
@@
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
@@
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
Since $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded, the first term on the right-hand side tends to zero due to (3.32). Since $1-\lambda_{l}\leq 1$, the second term on the right-hand side can be estimated by
@@
Due to (3.8) and Corollary 3.1, it follows that (3.34) tends to zero as $m\to\infty$. This means $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]\to 0$. Analogously, we have $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]\to 0$. Thus, result holds. ∎
In the following theorem, we establish the convergence of non-noisy iterates.
@@
Suppose that all the assumptions of Lemma 3.1 and (3.30) hold. Then, the IRSGD method (3.1) corresponding to $\delta=0$ generates a sequence of iterates $\{u_{k}\}_{k\geq 1}$ that converges to an element of $\mathcal{D}_{\rho+c(\rho)}(u^{\dagger})$ in $U$, i.e.,
$\lim_{k\to 0}\mathbb{E}[\|u_{k}-\hat{u}\|^{2}]=0,\quad\hat{u}\in\mathcal{D}_{\rho+c(\rho)}(u^{\dagger}).$
@@
The proof is similar to [2, Theorem 1]. ∎
3.4. Convergence for noisy case
The next result demonstrates the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$. Specifically, for any fixed integer $k\geq 0$ and any sample path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, it holds that $\|u_{k}^{\delta}-u_{k}\|^{2}\to 0$ as $\delta\to 0$. As a result, the expectation $\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]\to 0$ as $\delta\to 0$, confirming the pathwise stability of the iterates for each realization within the filtration $\mathcal{F}_{k}$.
@@ 3.4. Convergence for noisy case
Let all the assumptions of Lemma 3.1 be satisfied. For any fixed $k\in\mathbb{N}$ and any path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the IRSGD iterates along the path for exact data $y$ and noisy data $y^{\delta}$, respectively. Then
@@
We use induction to prove the assertion. For $k=0$ it is trivial. Suppose that it holds for all indices up to $k$ and any path in $\mathcal{F}_{k}$. From (3.1), for any fixed path $(i_{1},\ldots,i_{k})$, we have
@@
$+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
@@
Next, we claim that for any specific value of $k$, (i) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded, and (ii) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$ is uniformly bounded. Indeed, by (3.2) and (3.3), we derive that
@@
This, together with an induction argument, establishes the claim (i). Similarly, we note that
@@
and consequently 3.4 and 3.4 derive that
@@
This, along with mathematical induction establishes claim (ii). Let $c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|$. Then, it follows from 3.4, 3.4 and (3.2) that
@@
Finally, the result follows from the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, i.e.,
$\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0\implies\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
This completes the proof. ∎
The final result of this section establishes the regularizing nature of IRSGD method.
@@
Let all the assumptions of Lemma 3.1 hold. Let $\{y^{\delta}\}$ be the sequence of noisy data such that Assumption 3.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1. Then there exist a solution $\hat{u}$ of (1.2) such that
@@
Let $\{y^{\delta_{m}}\}$ be a subsequene of $\{y^{\delta}\}$, and let $K:=\lim\inf_{m\to\infty}k_{*}(y^{\delta_{m}}).$ By taking a subsequence of $\{y^{\delta_{m}}\}$ if necessary, we may assume that $K=\lim_{m\to\infty}k_{*}(y^{\delta_{m}})$ and, according to Lemma 3.5, we have
((3.37)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty.$
for all $0\leq k\leq K.$ Let $\hat{u}$ be the limit of $u_{k}$, which is a solution of (1.2) in $B_{\rho+c(\rho)}(u_{0}).$ First, we assume that $K<\infty,$ then we have $k_{*}(y^{\delta_{m}})=K$ for large $m.$ Thus, from Lemma 3.2 it follows that
$\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
By engaging (3.37) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the iteration terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
If $K=\infty,$ then Theorem 3.1 and Lemma 3.5 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
@@
At this stage, we apply the full conditional on (3.16) to arrive at
@@
where $C_{2}=2\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})$ and $C_{3}=\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that

### v3

@@ 2.2. Analysis for Heuristic Rule 1
((2.26)) $\Omega(1+\eta)\mathbb{E}[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
Hence, as a result, we get $\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2}.$ By incorporating Jensen’s inequality with $f(x)=x^{2}$, we conclude that
This concludes the proof. ∎
3. Convergence Analysis
In this section, we prove that $u^{\delta}_{k_{*}}$ converges almost surely to a solution of (1.1) as $\delta\to 0$. We first analyze the exact-data analogue of Algorithm 1 and establish its almost-sure convergence. Next, we show a stability property along each sample path that connects the noisy- and exact-data cases. These results together yield the main convergence theorem. The exact-data version of Algorithm 1 is given below.
Algorithm 3 Exact-data IRSGD
1:Given: $F_{i},y_{i}^{\dagger}$ for all $i\in\{0,1,\ldots,P-1\}$
2:Initialize: $u_{0}=u^{(0)}$
3:for $k\geq 0$ do
4:  Choose $\omega_{k}\in[\omega,\Omega]$ and the weight parameter $\lambda_{k}$ such that
$0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}.$
5:  Randomly pick an index $i_{k}$ from $\{0,1,\ldots,P-1\}$ uniformly.
6:  Update
$u_{k+1}=u_{k}-\omega_{k}F^{\prime}_{i_{k}}(u_{k})^{*}(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})-\lambda_{k}(u_{k}-u^{(0)}),$
where $F^{\prime}_{i_{k}}$ is the Fréchet derivative of $F_{i_{k}}$, and $u^{(0)}$ is the initial guess.
7:end for
[…]
((3.1)) $\daleth=\left\{\sum_{l=0}^{\infty}\sum_{i=0}^{P-1}\|F_{i}(u_{l})-y_{i}^{\dagger}\|^{2}<\infty\right\},$
and show that this event occurs almost surely, i.e., $\mathbb{P}(\daleth)=1$.
@@ 3. Convergence Analysis
Let the assumptions of Lemma 2.1 be satisfied. Additionally, assume that IRSGD method (1.11) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty,$
where $D_{1}$ is defined in (2.22). In particular, for the case $y^{\delta}=y^{\dagger}$, we have
$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
[…]
@@
$D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)\mathbb{E}[k_{*}\|y^{\delta}-y^{\dagger}\|^{2}].$
@@
Thus, result. The case $y^{\delta}=y^{\dagger}$ can be handled similarly. ∎
@@
If $\lambda_{\text{max}}=0$, the parameter $\kappa$ in conditions (2.6) and (2.22) can be set to zero. This adjustment mirrors the conditions required to establish the same results for SGD method (1.9).
3.1. Convergence for exact-data case
In this subsection, we establish the convergence of Algorithm 3. Let us assume that
((3.2)) $D_{2}:=(2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})>0.$
@@ 3.1. Convergence for exact-data case
Suppose all the conditions of Lemma 2.1 are satisfied. Additionally, assume that ((3.2)) holds. Then, the sequence of iterates $\{u_{k}\}_{k\geq 1}$ produced by Algorithm 3 is a Cauchy sequence almost surely.
@@
Firstly, we define $e_{k}:=u_{k}-u^{\dagger}$, for a solution $u^{\dagger}$ of (1.1). As shown in Lemma 2.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$. Let it converge to some $\epsilon\geq 0$. It follows directly from 2.2 that, for $\delta=0$ we have
@@
From the definition of $F$ and the measurability of $u_{k}^{\delta}$ concerning $\mathcal{F}_{k}$, we deduce that
@@
$\quad+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
$\quad-\frac{1}{P}\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
Taking the total conditional and using (3.2), we get
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+\frac{D_{2}}{P}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
@@
for $m<k,$ where we have used the following result
@@
These bounds, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, directly imply
@@
Note that the operator $\mathbb{E}[\langle\cdot,\cdot\rangle]$ defines an inner product. Now, for any $n\geq k$, we pick an index $m$ such that $k\leq m\leq n$ and
((3.5)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\text{for all }k\leq i\leq n.$
Using the triangular inequality
((3.6)) $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]^{1/2},$
along with the identities
((3.7)) $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
((3.8)) $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
we aim to prove the desired result. To that end, it suffices to show that both $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ converge to zero as $k\to\infty$.
Since $\mathbb{E}[\|e_{k}\|^{2}]\to\epsilon$ as $k\to\infty$, the final two terms on the right-hand sides of the above identities vanish in the limit, i.e., $\epsilon-\epsilon=0$. Thus, it remains to show that $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also tends to zero as $k\to\infty$. The same argument applies to $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]$.
To show this, we recall the recursive structure of $u_{k}$, from which we obtain
$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}),$
for $m<k$. This formula allows us to determine that
@@
$\quad+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
@@
$\quad+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
Due to (3.4), the first term on the right-hand side approaches zero because $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded. In addition to that, by using the inequality $1-\lambda_{k}\leq 1$ along with (2.7), one can estimate the second term on the right-hand side as
@@
Now from Corollary 3.1, it follows that (3.9) approaches zero as $k\to\infty$. This means $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]\to 0$. Analogously, we have $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]\to 0$. Hence from ((3.7)) and ((3.8)) we have
$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]\to 0\quad\text{and}\quad\mathbb{E}[\|e_{m}-e_{k}\|^{2}]\to 0$
as $k\to\infty.$ Hence, the desired result follows directly from inequality ((3.6)). ∎
[…]
@@
[…]
@@
Lemma 3.1 establishes that the sequence of iterates $\{u_{k}\}_{k\geq 1}$ produced by Algorithm 3 forms a Cauchy sequence in $U$. Consequently, the sequence admits a limit, denoted by $u^{\dagger}\in U$. Since the operator $F$ is continuous, it follows that $F(u_{k})\to F(u^{\dagger})$ as $k\to\infty$. Moreover, Corollary 3.1 guarantees that
$\lim_{k\to\infty}\mathbb{E}\left[\|F(u_{k})-y^{\dagger}\|^{2}\right]=0,$
[…]
3.2. Regularization Quality
The primary goal of this subsection is to demonstrate that the IRSGD method possesses the regularizing property when employed with the proposed heuristic stopping rule. Our approach begins by establishing the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$, thereby linking Algorithm 1 for noisy data to its exact-data counterpart in Algorithm 3. The proof structure is adapted from the methodology outlined in [16, Lemma 3.6].
@@ 3.2. Regularization Quality
Assume that every condition in Lemma 2.1 is met. For any given $k\in\mathbb{N}$ and any realization of the sequence $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the iterates produced by the IRSGD method (1.11) for the exact and noisy data, respectively. Then
@@
We use induction to prove the assertion. For $k=0$ it is trivial. Assume that it is true for any path in $\mathcal{F}_{k}$ and for all indices up to $k$. According to (1.11), for any fixed path $(i_{1},\ldots,i_{k})$, we get
@@
$\quad+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
@@
We proceed to establish the following statements for any fixed $k$
  1. (i)
    The supremum $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded.
  2. (ii)
    The supremum of the difference $\|u_{k}^{\delta}-u_{k}\|$ over all realizations in $\mathcal{F}_{k}$, i.e., $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$, is uniformly bounded.
To prove claim (i), we invoke ((2.1)) and ((2.2)), which yields that
@@
This recursive inequality implies that the sequence $\{\|u_{k}-\hat{u}\|\}$ grows at most linearly, and by applying an induction argument, we conclude that claim (i) is established. Similarly, to establish claim (ii), we note that
@@
and consequently 3.2 and 3.2 derive that
@@
This, together with a standard induction argument, confirms claim (ii). Consequently, leveraging both claims, we may now proceed to let
$c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|.$
Then, it follows from 3.2, 3.2 and (2.1) that
@@
Finally, by the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, we obtain $\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0,$ which, in turn, implies $\lim_{\delta\to 0^{+}}\mathbb{E}\big{[}\|u_{k}^{\delta}-u_{k}\|^{2}\big{]}=0.$ This completes the proof. ∎
We are now prepared to formally demonstrate the central result of this subsection, which concerns the regularization property of the IRSGD method.
@@
Assume that every condition in Lemma 2.1 is met. Suppose that $\{y^{\delta}\}$ is the noisy data sequence such that Assumption 2.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1. Then there exist a solution $\hat{u}$ of (1.2) such that
@@
[…]
$\lim_{\delta\to 0}\|u_{k_{*}(y^{\delta})}^{\delta}-\hat{u}\|^{2}]=0\quad\text{on the event}\daleth.$
[…]
((3.12)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty\ \ \text{for all}\ 0\leq k\leq K.$
There are two possible cases to consider:
Case I: Suppose $K<\infty$. Then, for sufficiently large $m$, we have $k_{*}(y^{\delta_{m}})=K$. Thus, from Lemma 2.2 it follows that
$\mathbb{E}\left[\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\right]\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
By engaging (3.12) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the method terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
Case II: Suppose $K=\infty$. Then Theorem 3.1 and Lemma 3.2 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
@@
At this stage, we apply the full conditional on 2.2 to arrive at
@@
where $C_{2}=\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)$ and $C_{3}=\frac{M\Omega(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that